How many liters of chromium are needed to react with an excess of CuSO4 to produce 7.50 moles Cu?

1
Expert's answer
2021-03-03T06:30:22-0500

Cr + CuSO4 = CrSO4 + Cu

According to the equation

n(Cu) = n(Cr) = 7.5 moles

V(Cr) = n*Vm = 7.5mol*22.4L/mol = 168 L